# Relatively big projective modules and their applications to direct sum decompositions

### v1

@@ 4.3. An application to direct sums of finitely generated modules
To transfer the information about projective modules over the endomorphism ring of a finitely generated module $M$ to $\mbox{\rm Add}(M)$, we shall use the following well-known result.
@@
([libro, Theorem 4.7]) Let $R$ be a ring, and let $M$ be a right $R$-module with endomorphism ring $\Lambda=\mathrm{End}_{R}(M)$. Then the functor $\mathrm{Hom}_{R}(M,-)$ induces a category equivalence between $\mbox{\rm add}(M)$ and the category of finitely generated projective right $\Lambda$-modules.
Assume, in addition, that $M_{R}$ is finitely generated. Then the functor $\mathrm{Hom}_{R}(M,-)$ induces a category equivalence between $\mbox{\rm Add}(M)$ and the category of projective right $\Lambda$-modules.
Remark 4.12.
In the notation of Proposition 4.11, the inverse of the equivalence induced by $\mathrm{Hom}_{R}(M,-)$ is given by the functor $-\otimes_{\Lambda}M$.
@@
Proof.  As $\mathrm{End}_{\Lambda}(M)$ is a module-finite $R$-algebra, we can apply Corollary 4.10, to deduce that all countably generated projective modules over $\mathrm{End}_{\Lambda}(M)$ that are not finitely generated are relatively big. Hence, they cannot be indecomposable. The conclusion follows from Proposition 4.11.
@@
In general, if $\Lambda$ is a PI-ring, then the endomorphism ring of a finitely generated right $\Lambda$-module is also a PI-ring, see the argument of [PS, Theorem 2].
@@ 5. Projective versus locally projective modules
[AHP, Lemma 2.8] Let $R$ be a domain, and let $\Lambda$ be an $R$-algebra. Let $\Sigma$ be a multiplicative subset of $R$. If $M$ is a finitely generated right $\Lambda$-module that is torsion-free as an $R$-module, then the canonical injective homomorphism
@@
To prove $(iii)$, let $x\in\operatorname{Tr}_{\Lambda\otimes_{R}R_{\Sigma}}(M\otimes_{R}R_{\Sigma})$. Then there is an element $t\in\Sigma$, and $f_{1},\dotsc,f_{n}\in\operatorname{Hom}_{\Lambda\otimes_{R}R_{\Sigma}}(M\otimes_{R}R_{\Sigma},\Lambda\otimes_{R}R_{\Sigma})$, $m_{1},\dotsc,m_{n}\in M$ such that $x=\sum_{i=1}^{n}f_{i}\left(m_{i}\otimes\frac{1}{t}\right)$. Let
@@
denote the canonical homomorphism. By Lemma 5.2, $\varphi$ is an isomorphism, so there is an element $r\in\Sigma$ and $g_{1},\dotsc,g_{n}\in\operatorname{Hom}_{\Lambda}(M,\Lambda)$ such that $\varphi(g_{i}\otimes\frac{1}{r})=f_{i}$, for each $i=1,\dotsc,n$. Then
@@
The following result is necessary to give a version of the Package Deal Theorem for projective modules.
@@
The domain $R$ has finite character if $R/I$ is a semilocal ring for every non-zero ideal $I$ of $R$. If $R$ is $h$-local then it also satisfies that $R/\mathfrak{p}$ is a local domain for every non-zero prime ideal $\mathfrak{p}$ of $R$. Equivalently, $R$ is an $h$-local domain if and only if $R/I$ is semiperfect for every non-zero ideal $I$ of $R$. For characterizations of $h$-local domains the reader is referred to [olberding].
@@
Let $R$ be a domain, and let $\Lambda$ be an $R$-algebra. Let $\Sigma$ be a multiplicative subset of $R$. Let $U$ be a right $\Lambda$-module which is torsion-free as an $R$-module. If $V$ is a finitely generated $\Lambda$-submodule that is torsion-free and such that $V_{\Sigma}$ is isomorphic to a submodule of $U_{\Sigma}$, then $U$ has a submodule isomorphic to $V$.
Proof.  By hypothesis, and because $V$ is torsion-free, $U_{\Sigma}$ has a submodule $V^{\prime}$ isomorphic to $V$. Since $V^{\prime}$ is finitely generated, there exists $d\in\Sigma$ such that $V^{\prime}\subseteq\frac{1}{d}U\subseteq U_{\Sigma}$. Then $dV^{\prime}\leq U$ is the module we are looking for.
@@
In the following result, we specialize to two-sided ideals of $\Lambda$ and assume that $\Lambda_{Q}$ is simple artinian.
@@
    there exists $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$ maximal ideals of $R$ such that $\Lambda/I=\prod_{i=1}^{\ell}\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$.
@@
The following results, which are known as Package Deal Theorems, were proved in [levyodenthal2] for the case of algebras over noetherian rings of Krull dimension 1. The name of these theorems comes from the idea of having a “package” of modules over a localization and finding a module over the non-localized algebra whose localizations are those given in the “package”. In [AHP] Package Deal Theorems were given for localizations of submodules and localizations of trace ideals. We recall here the statements.
@@
Proof.  $(ii)\Rightarrow(i)$. Let $\mathcal{M}$ be the finite set of maximal ideals of $R$ such that $P(\mathfrak{m})\not\cong\Lambda_{\mathfrak{m}}^{r}$. Consider a module $F=\Lambda^{r}$. Since $\Lambda$ is torsion-free as an $R$-module, for every $\mathfrak{m}\in\mathcal{M}$ there exists a submodule $F(\mathfrak{m})$ of $F_{\mathfrak{m}}$ such that $F(\mathfrak{m})\cong P(\mathfrak{m})$. For $\mathfrak{n}\in\operatorname{mSpec}R\setminus\mathcal{M}$ set $F(\mathfrak{n})=F_{\mathfrak{n}}$. By Package Deal Theorem 5.9, there is a finitely generated right $\Lambda$-module $P$ contained in $F$ such that $P_{\mathfrak{m}}=F(\mathfrak{m})\cong P(\mathfrak{m})$ for every $\mathfrak{m}\in\mathcal{\operatorname{mSpec}}R$. By Corollary 5.4, $P$ is a projective module. Also note that $\operatorname{Tr}_{\Lambda}(P)_{\mathfrak{m}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(P_{\mathfrak{m}})$ by Lemma 5.3.
@@ 6. Semiperfect rings and locally semiperfect rings
We start this section by examining quickly how the monoids of projective modules we have introduced are for the case of semiperfect rings. We follow the notation and conventions of [andersonfuller, §27].
@@ 6.1. Projective modules over semiperfect rings and their trace ideals
Let $\Lambda$ be a semiperfect ring, and let $I$ be an ideal of $\Lambda$. Let $P^{\prime}$ be a finitely generated, projective right $\Lambda/I$-module. Then there exists a finitely generated projective right $\Lambda$-module $X$ such that $X/XI\cong P^{\prime}$.
@@
    Let $P$ be a projective right $\Lambda$-module with trace $I$. Assume that $I/IJ(\Lambda)\cong\sum_{i\in A}\left((e_{i}+J(\Lambda))\Lambda/J(\Lambda)\right)^{\ell_{i}}$ for suitable $\ell_{i}>0$ and $A\in\mathcal{P}(\{1,\dots,t\})$. Then $P\cong\bigoplus_{i\in A}\left(e_{i}\Lambda\right)^{(L_{i})}$, for suitable non-empty sets, $L_{i}$.
@@
    Keeping the notation from $(ii)$ and $(iii)$, $\Lambda^{\prime}:=\Lambda/I$ is a semiperfect ring and $e_{i}^{\prime}:=e_{i}+I,i\not\in A$ is a basic set of idempotents of $\Lambda/I$. Moreover, let $P^{\prime}\cong\bigoplus_{i\not\in A}(e_{i}^{\prime}\Lambda^{\prime})^{x_{i}}$ be a finitely generated projective $\Lambda^{\prime}$-module and let $P$ be a finitely generated projective $\Lambda$-module such that $P/PI\cong P^{\prime}$. Then there are non-negative integers $s_{i},i\not\in A$ such that $P\cong\bigoplus_{i\not\in A}(e_{i}\Lambda)^{x_{i}}\oplus\bigoplus_{i\in A}(e_{i}\Lambda)^{s_{i}}$.
Proof.  The proof of the four statements is more or less routine. One just has to take into account that any projective right (left) $\Lambda$-module is isomorphic to a unique direct sum of the indecomposable projective modules $P_{i}=e_{i}\Lambda$ ($Q_{i}=\Lambda e_{i}$), for $i=1,\dots,t$ (see for example [libro, Theorem 3.10]). Moreover $\mathrm{Tr}_{\Lambda}(P_{i})=\mathrm{Tr}_{\Lambda}(Q_{i})=\Lambda e_{i}\Lambda$.
@@
    The canonical isomorphism between $V_{r}(\Lambda):=V(\Lambda)\cong{\mathbb{N}}_{0}^{t}$ and $V_{\ell}(\Lambda)$ extends to an isomorphism $V_{r}^{*}(\Lambda)\cong V_{\ell}^{*}(\Lambda):=V^{*}(\Lambda)\cong({\mathbb{N}}_{0}^{*})^{t}$;
@@
    Let $P$ be a projective right $\Lambda$-module. Then there exists a two-sided ideal $I$ of $\Lambda$ minimal with respect to the property $P/PI$ is finitely generated. Moreover, $I$ is the trace of a finitely generated projective module $Q$ and $P$ is $I$-big.
@@
    $V^{*}(\Lambda)=V(\Lambda)\sqcup B(\Lambda)$.
Proof.  $(i)$ Let $e_{1},\dots,e_{t}$ be a basic set of idempotents of $\Lambda$. Then $P_{i}=e_{i}\Lambda$ and $Q_{i}=\Lambda e_{i}\cong\mathrm{Hom}_{\Lambda}(P_{i},\Lambda)$, for $i=1,\dots,t$, are representatives of the isomorphism classes of indecomposable projective right and left $\Lambda$-modules. As any projective module is a direct sum of indecomposables, in a unique way up to isomorphism, the first statement is clear.
$(ii)$ If $P$ is a projective right $\Lambda$-module, then $\{1,\dots,t\}=A\cup B$, where $A$ and $B$ are disjoint, and $P\cong\left(\bigoplus_{i\in A}P_{i}^{(A_{i})}\right)\oplus\left(\bigoplus_{i\in B}P_{i}^{(B_{i})}\right)$, where $A_{i}$ is a finite set (possibly empty) for any $i\in A$ and $B_{i}$ is an infinite set for any $i\in B$. Then
@@
$(iii)$ By Proposition 6.2, $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$ so we can apply Corollary 2.12 to deduce that $B_{r}(\Lambda):=B(\Lambda)$ is canonically isomorphic to $B_{\ell}(\Lambda)$. The statement $(ii)$ implies that $V^{*}(\Lambda)=V(\Lambda)\sqcup B(\Lambda)$.

### v2

@@ 4.3. An application to direct sums of finitely generated modules
Let $M$ be a right $\Lambda$-module. By $\mbox{\rm add}(M)$ we mean the full subcategory of $\Lambda$-modules that are isomorphic to a direct summand of $M^{n}$ for some natural number $n$. By $\mbox{\rm Add}(M)$ we denote the full subcategory of $\Lambda$-modules that are isomorphic to a direct summand of an arbitrary direct sum of copies of $M$. It is also convenient to consider the subcategory $\mbox{\rm Add}_{\aleph_{0}}(M)$ of direct summands of $M^{(\omega)}$.
Note that $\mbox{\rm Add}(\Lambda)$ ($\mbox{\rm add}(\Lambda)$) is the category of (finitely generated) projective right $\Lambda$-modules. While $\mbox{\rm Add}_{\aleph_{0}}(\Lambda)$ is the category of countably generated projective right $\Lambda$-modules.
It is possible to transfer the information about projective modules over the endomorphism ring of a finitely generated module $M$ to $\mbox{\rm Add}(M)$, via the following well-known result.
@@
[libro, Theorem 4.7] Let $\Lambda$ be a ring, and let $M$ be a right $\Lambda$-module. Let $S=\operatorname{End}_{\Lambda}(M)$. Then the functors $\operatorname{Hom}_{\Lambda}(M,-)$ and $\otimes_{S}M$ induce an equivalence between $\mbox{\rm add}(M)$ and $\mbox{\rm add}(S)$.
Moreover, if $M$ is finitely generated, then they induce an equivalence between $\mbox{\rm Add}(M)$ and $\mbox{\rm Add}(S)$. This equivalence restricts to an equivalence between $\mbox{\rm Add}_{\aleph_{0}}(M)$ and $\mbox{\rm Add}_{\aleph_{0}}(S)$.
@@
Proof. As $\mathrm{End}_{\Lambda}(M)$ is a module-finite $R$-algebra, we can apply Corollary 4.10 to deduce that all countably generated projective modules over $\mathrm{End}_{\Lambda}(M)$ are relatively big. Hence, by Remark 1.18, the indecomposable projectives over $\operatorname{End}_{\Lambda}(M)$ are finitely generated. The conclusion follows from Proposition 4.11.
@@
In general, if $\Lambda$ is a PI-ring, then the endomorphism ring of a finitely generated right $\Lambda$-module is also a PI-ring (see the argument of [PS, Theorem 2]).
@@ 5. Projective versus locally projective modules
[AHP, Lemma 2.8] Let $R$ be a commutative ring, and let $\Lambda$ be an $R$-algebra. Let $\Sigma$ be a multiplicative subset of $R$ consisting of non-zero divisors. If $M$ is a finitely generated right $\Lambda$-module that is torsion-free as an $R$-module, then the canonical injective homomorphism
@@
To prove $(iii)$, let $x\in\operatorname{Tr}_{\Lambda\otimes_{R}R_{\Sigma}}(M\otimes_{R}R_{\Sigma})$. Then there is an element $t\in\Sigma$, and elements $f_{1},\dotsc,f_{n}\in\operatorname{Hom}_{\Lambda\otimes_{R}R_{\Sigma}}(M\otimes_{R}R_{\Sigma},\Lambda\otimes_{R}R_{\Sigma})$ and $m_{1},\dotsc,m_{n}\in M$ such that $x=\sum_{i=1}^{n}f_{i}\left(m_{i}\otimes\frac{1}{t}\right)$. Let
@@
denote the canonical homomorphism. By Lemma 5.2, $\varphi$ is an isomorphism, so there is an element $r\in\Sigma$ and elements $g_{1},\dotsc,g_{n}\in\operatorname{Hom}_{\Lambda}(M,\Lambda)$ such that $\varphi(g_{i}\otimes\frac{1}{r})=f_{i}$ for each $i=1,\dotsc,n$. Then
@@
It is easy to prove that the notion of relatively big projective module behaves well with respect to localization.
Corollary 5.4.
Let $\Lambda$ be an algebra over a commutative ring $R$. Let $J\in\mathcal{T}(\Lambda)$. Let $\Sigma$ be a multiplicatively closed subset of $R$ consisting of non-zero divisors. If $P$ is a relatively $J$-big projective right $\Lambda$-module, then $P_{\Sigma}$ is a relatively $J_{\Sigma}$-big projective right $\Lambda_{\Sigma}$-module.
Proof. By Lemma 5.3, $J_{\Sigma}\in\mathcal{T}(\Lambda_{\Sigma})$. Since $P/PJ$ is finitely generated, so is $(P/PJ)_{\Sigma}\cong P_{\Sigma}/P_{\Sigma}J_{\Sigma}$.
Let $P^{\prime}$ be a countably generated projective right $\Lambda$-module with trace $J$. Since $P\cong P\oplus Q^{(\omega)}$ as $\Lambda$-modules, also $P_{\Sigma}\cong P_{\Sigma}\oplus(P^{\prime})_{\Sigma}^{(\omega)}$ as $\Lambda_{\Sigma}$-modules. This finishes the proof of the claim.
The following result is necessary to provide a version of the Package Deal Theorem for projective modules.
@@
The domain $R$ has finite character if $R/I$ is a semilocal ring for every non-zero ideal $I$ of $R$. If $R$ is $h$-local, then it also satisfies that $R/\mathfrak{p}$ is a local domain for every non-zero prime ideal $\mathfrak{p}$ of $R$. Equivalently, $R$ is an $h$-local domain if and only if $R/I$ is semiperfect for every non-zero ideal $I$ of $R$. For characterizations of $h$-local domains, the reader is referred to [olberding].
@@
Let $R$ be a domain, and let $\Lambda$ be an $R$-algebra. Let $\Sigma$ be a multiplicative subset of $R$. Let $U$ be a right $\Lambda$-module which is torsion-free as an $R$-module. If $V$ is a finitely generated $\Lambda$-module that is torsion-free and such that $V_{\Sigma}$ is isomorphic to a submodule of $U_{\Sigma}$, then $U$ has a submodule isomorphic to $V$.
Proof. By hypothesis, and because $V$ is torsion-free, $U_{\Sigma}$ has a submodule $V^{\prime}$ that is isomorphic to $V$. Since $V^{\prime}$ is finitely generated, there exists $d\in\Sigma$ such that $V^{\prime}\subseteq\frac{1}{d}U\subseteq U_{\Sigma}$. Then $dV^{\prime}\leq U$ is the module we are looking for.
@@
In the following result, we specialize in two-sided ideals of $\Lambda$ and assume that $\Lambda_{Q}$ is a simple artinian ring.
@@
    There exist $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$ maximal ideals of $R$ such that $\Lambda/I=\prod_{i=1}^{\ell}\Lambda_{\mathfrak{m}_{i}}/I_{\mathfrak{m}_{i}}$.
@@
The following results, which are known as Package Deal Theorems, were proved in [levyodenthal2] for the case of algebras over noetherian rings of Krull dimension 1. The name of these theorems comes from the idea of having a “package” of modules over a localization and finding a module over the non-localized algebra whose localizations are those given in the “package”. In [AHP] Package Deal Theorems were presented for the localizations of submodules and the localizations of trace ideals. We recall the statements here.
@@
Proof. $(ii)\Rightarrow(i)$. Let $\mathcal{M}$ be the finite set of maximal ideals of $R$ such that $P(\mathfrak{m})\not\cong\Lambda_{\mathfrak{m}}^{r}$. Consider a module $F=\Lambda^{r}$. Since $\Lambda$ is torsion-free as an $R$-module, for every $\mathfrak{m}\in\mathcal{M}$ there exists a submodule $F(\mathfrak{m})$ of $F_{\mathfrak{m}}$ such that $F(\mathfrak{m})\cong P(\mathfrak{m})$. For $\mathfrak{n}\in\operatorname{mSpec}(R)\setminus\mathcal{M}$ set $F(\mathfrak{n})=F_{\mathfrak{n}}$. By Package Deal Theorem 5.10, there is a finitely generated right $\Lambda$-module $P$ contained in $F$ such that $P_{\mathfrak{m}}=F(\mathfrak{m})\cong P(\mathfrak{m})$ for every $\mathfrak{m}\in\mathcal{\operatorname{mSpec}}(R)$. By Corollary 5.5, $P$ is a projective module. Also note that $\operatorname{Tr}_{\Lambda}(P)_{\mathfrak{m}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(P_{\mathfrak{m}})$ by Lemma 5.3.
@@ 6. Semiperfect rings and locally semiperfect rings
We start this section by describing the monoids of projective modules for semiperfect rings. We follow the notation and conventions of [andersonfuller, §27].
@@ 6.1. Projective modules over semiperfect rings and their trace ideals
Let $\Lambda$ be a semiperfect ring, and let $I$ be an ideal of $\Lambda$. Let $P^{\prime}$ be a finitely generated projective right $\Lambda/I$-module. Then, there exists a finitely generated projective right $\Lambda$-module $X$ such that $X/XI\cong P^{\prime}$.
@@
    Let $P$ be a projective right $\Lambda$-module with trace ideal $I$. Assume that $I/(I\cap J(\Lambda))\cong\sum_{i\in A}\left((e_{i}+J(\Lambda))\Lambda/J(\Lambda)\right)^{\ell_{i}}$ for suitable $\ell_{i}>0$ and $A\in\mathcal{P}(\{1,\dots,t\})$. Then $P\cong\bigoplus_{i\in A}\left(e_{i}\Lambda\right)^{(L_{i})}$, for suitable non-empty sets, $L_{i}$.
@@
    Keeping the notation from $(ii)$ and $(iii)$, $\Lambda^{\prime}:=\Lambda/I$ is a semiperfect ring and $e_{i}^{\prime}:=e_{i}+I,i\not\in A$ is a basic set of idempotents of $\Lambda/I$. Moreover, let $P^{\prime}\cong\bigoplus_{i\not\in A}(e_{i}^{\prime}\Lambda^{\prime})^{x_{i}}$ be a finitely generated projective $\Lambda^{\prime}$-module, and let $P$ be a finitely generated projective $\Lambda$-module such that $P/PI\cong P^{\prime}$. Then there are non-negative integers $s_{i},i\not\in A$ such that $P\cong\bigoplus_{i\not\in A}(e_{i}\Lambda)^{x_{i}}\oplus\bigoplus_{i\in A}(e_{i}\Lambda)^{s_{i}}$.
Proof. The proof of the four statements is more or less routine. One just has to take into account that any projective right (left) $\Lambda$-module is isomorphic to a unique direct sum of the indecomposable projective modules $P_{i}=e_{i}\Lambda$ ($Q_{i}=\Lambda e_{i}$), for $i=1,\dots,t$ (see, for example, [libro, Theorem 3.10]). Moreover $\mathrm{Tr}_{\Lambda}(P_{i})=\mathrm{Tr}_{\Lambda}(Q_{i})=\Lambda e_{i}\Lambda$.
To give explicit descriptions of the monoids of projective modules over certain classes of rings, we introduce the monoid/semiring ${\mathbb{N}}_{0}^{*}$.
Notation 6.3.
Let $\mathbb{N}_{0}=\{0,1,2,\dots\}$, and let $\mathbb{N}=\{1,2,\dots\}$. We denote by ${\mathbb{N}}_{0}^{*}=\mathbb{N}_{0}\cup\{\infty\}$ the monoid with the sum extending the one of ${\mathbb{N}}_{0}$ and with the additional rule that $x+\infty=\infty+x=\infty$ for any $x\in{\mathbb{N}}_{0}^{*}$. We also see ${\mathbb{N}}_{0}^{*}$ as a semiring with the product extending the one of ${\mathbb{N}}_{0}$ together with the rules $x\cdot\infty=\infty\cdot x=\infty$ for any $x\in{\mathbb{N}}_{0}^{*}\setminus\{0\}$, and $0\cdot\infty=\infty\cdot 0=0$.
@@
    The canonical isomorphism between $V(\Lambda):=V_{r}(\Lambda)\cong{\mathbb{N}}_{0}^{t}$ and $V_{\ell}(\Lambda)$ extends to an isomorphism between $V^{*}(\Lambda):=V_{r}^{*}(\Lambda)\cong({\mathbb{N}}_{0}^{*})^{t}$ and $V_{\ell}^{*}(\Lambda)$.
@@
    Let $P$ be a projective right $\Lambda$-module. Then there exists a two-sided ideal $I$ of $\Lambda$ that is minimal with respect to the property that $P/PI$ is finitely generated. Moreover, $I$ is the trace of a finitely generated projective right module.
@@
    Every countably generated projective module is relatively big, so that $V^{*}(\Lambda)=V(\Lambda)\sqcup B(\Lambda)$.
Proof. Let $e_{1},\dots,e_{t}$ be a basic set of idempotents of $\Lambda$. Then $P_{i}=e_{i}\Lambda$ and $Q_{i}=\Lambda e_{i}\cong\mathrm{Hom}_{\Lambda}(P_{i},\Lambda)$, for $i=1,\dots,t$, are representatives of the isomorphism classes of indecomposable projective right and left $\Lambda$-modules. Recall that any projective right or left module over $\Lambda$ is a direct sum of indecomposables, uniquely up to isomorphism.
$(i)$ Let $P$ be a countably generated projective right $\Lambda$-module. Then there exist $A_{1},\dots,A_{t}$ countable sets such that $P\cong\bigoplus_{i=1}^{t}P_{i}^{(A_{i})}$.
Then the assignment
$P]=\left[\bigoplus_{i=1}^{t}P_{i}^{(A_{i})}\right]\mapsto\left[\bigoplus_{i=1}^{t}Q_{i}^{(A_{i})}\right]$
gives an isomorphism of monoids $V_{r}^{*}(\Lambda)\to V_{\ell}^{*}(\Lambda)$ that extends the canonical isomorphism $V_{r}(\Lambda)\to V_{\ell}(\Lambda)$ from Remark 2.6.
Moreover, the assignment $P]\mapsto(n_{1},\dots,n_{t})$, where for $i=1,\dots,t$
$n_{i}=\begin{cases}|A_{i}|&\text{if $A_{i}$ is finite}\\ \infty&\text{otherwise},\end{cases}$
induces an isomorphism $V_{r}^{*}(\Lambda)\to({\mathbb{N}}_{0}^{*})^{t}$ that restricts to an isomorphism $V_{r}(\Lambda)\to{\mathbb{N}}_{0}^{t}$.
$(ii)$ If $P$ is a projective right $\Lambda$-module, then $\{1,\dots,t\}=A\cup B$, where $A$ and $B$ are disjoint, and $P\cong\left(\bigoplus_{i\in A}P_{i}^{(A_{i})}\right)\oplus\left(\bigoplus_{i\in B}P_{i}^{(B_{i})}\right)$, where $A_{i}$ is a finite set (possibly empty) for any $i\in A$, and $B_{i}$ is an infinite set for any $i\in B$. Then, by Lemma 1.5,
@@
$(iii)$ By Proposition 6.2, $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$, so we can apply Corollary 2.9 to deduce that $B_{r}(\Lambda):=B(\Lambda)$ is canonically isomorphic to $B_{\ell}(\Lambda)$.
Statement $(ii)$ implies that any countably generated projective module is relatively big. Hence, $V^{*}(\Lambda)=V(\Lambda)\sqcup B(\Lambda)$.